# thanks: Corresponding author

### v1

@@ Abstract
We study the exotic $\Delta^{0}\Delta^{0}$ dibaryon state with quantum numbers $J^{P}=2^{+}$ within the framework of QCD sum rules. A tensor interpolating current is constructed to determine the mass and residue of this state, including explicit calculations of contributions from quark, gluon, and mixed vacuum condensates up to dimension nine. The numerical analysis yields a mass of $m_{\Delta^{0}\Delta^{0}}=2426^{+101}_{-108}~\mathrm{MeV}$ and a residue of $f_{\Delta^{0}\Delta^{0}}=\left(2.30^{+0.49}_{-0.45}\right)\times 10^{-4}~\mathrm{GeV}^{8}$. The predicted mass lies about $38~\mathrm{MeV}$ below the $2m_{\Delta^{0}}$ threshold, which indicates the possible existence of a bound molecular configuration. Consequently, the tensor $\Delta^{0}\Delta^{0}$ dibaryon may serve as a promising candidate for future experimental searches.
@@ III Numerical analysis
Here, we impose the following criteria: The upper bound of the Borel parameter $M^{2}$ is determined by ensuring that the pole contribution is considerably large compared to each of the higher excited states and continuum. In technical language, we require that the pole contribution be as maximum as possible in the exotic dibaryon channel: We set $\mathrm{PC}\geq 0.2$. Moreover, the lower bound of $M^{2}$ is determined by the requirement of convergence of the OPE. This criterion implies that the perturbative contribution must dominate over the non-perturbative terms, and the effects of higher-dimensional operators should remain suppressed. To ensure this, we require that the combined contribution of the last three non-perturbative terms does not exceed $5\%$ of the total (perturbative plus non-perturbative) contributions, i.e., $R(M^{2})\leq 0.05$. Moreover, the chosen range for $s_{0}$ should lie significantly above the production threshold of two $\Delta^{0}$ baryons, which is approximately $4m_{\Delta^{0}}^{2}\simeq 6.07~\mathrm{GeV}^{2}$. The range of $s_{0}$ may be selected as the minimal value that ensures a stable and reliable Borel window. Alternatively, $s_{0}$ can be fixed by fulfilling additional constraints so as to maximize the pole contribution (PC). Note that for multiquark hadrons, or more precisely exotic states, the available information is often scarce. In such situations, $s_{0}$ can be estimated based on the aforementioned constraints.
@@
It is important to emphasize that the present analysis excludes the influence of radiative corrections. Although the mass, extracted from the ratio of two sum rules, is relatively insensitive to such corrections, the residue calculated from an individual sum rule can exhibit considerable fluctuations. Furthermore, next-to-leading order (NLO) perturbative corrections to baryon sum rules have been reported as significant Krasnikov:1982ea ; Groote:1999zp . Similarly, while NLO $\alpha_{s}$ corrections to the perturbative component in the operator product expansion (OPE) for light tetraquark currents are sizable, their overall impact on the sum rules remains limited owing to the predominance of nonperturbative condensate contributions Groote:2014pva . Furthermore, given that the analysis of the $\Delta^{0}\Delta^{0}$ dibaryon with $J^{P}=2^{+}$ suggests the existence of a bound state, the estimation of its size can be interesting. Based on the definition of binding energy, the size of the $\Delta^{0}\Delta^{0}$ dibaryon can be determined using the following expression:
@@
where the binding energy is $B_{E}=2m_{\Delta^{0}}-m_{\Delta^{0}\Delta^{0}}$. In this expression, $r$ represents the distance between the components, while $\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}}$ corresponds to the reduced mass of the two-hadron system. The size is found to be approximately $r\simeq 0.90~\mathrm{fm}$, which is nearly twice that of the $\Delta^{0}\Delta^{0}$ dibaryon in the $J^{P}=0^{+}$ state, though both remain below the typical confinement radius of $1~\mathrm{fm}$ Guo:2017jvc . In contrast to the $\Delta^{0}\Delta^{0}$ dibaryon, which represents a light-quark exotic dibaryon, it is of particular interest to estimate the radius of the $\Omega\Omega$ dibaryon mentioned above. Based on the binding energy calculated in Ref. Chen:2019vdh , and assuming the same quantum state as considered in our study, the resulting radii are found to be approximately $r\simeq 1.24~\mathrm{fm}$ for the $J^{P}=0^{+}$ state and $r\simeq 0.97~\mathrm{fm}$ for the $J^{P}=2^{+}$ state. These results indicate that the tensor state remains within the typical confinement radius.

### v2

@@ Abstract
We study the exotic $\Delta^{0}\Delta^{0}$ dibaryon state with quantum numbers $J^{P}=2^{+}$ within the framework of QCD sum rules. A tensor interpolating current is constructed to determine the mass and residue of this state, including explicit calculations of contributions from quark, gluon, and mixed vacuum condensates up to dimension nine. The numerical analysis yields a mass of $m_{\Delta^{0}\Delta^{0}}=2426^{+101}_{-108}~\mathrm{MeV}$ and a residue of $f_{\Delta^{0}\Delta^{0}}=\left(2.30^{+0.49}_{-0.45}\right)\times 10^{-4}~\mathrm{GeV}^{8}$. The predicted mass lies about $38~\mathrm{MeV}$ below the $2m_{\Delta^{0}}$ threshold, which indicates the possible existence of a bound configuration. Consequently, the tensor $\Delta^{0}\Delta^{0}$ dibaryon may serve as a promising candidate for future experimental searches.
@@ III Numerical analysis
Here, we impose the following criteria: The upper bound of the Borel parameter $M^{2}$ is determined by ensuring that the pole contribution is considerably large compared to each of the higher excited states and continuum. In technical language, we require that the pole contribution be as maximum as possible in the exotic dibaryon channel: We set $\mathrm{PC}\geq 0.2$. Moreover, the lower bound of $M^{2}$ is determined by the requirement of convergence of the OPE. This criterion implies that the perturbative contribution must dominate over the non-perturbative terms, and the effects of higher-dimensional operators should remain suppressed. To ensure this, we require that the combined contribution of the last three non-perturbative terms does not exceed $5\%$ of the total (perturbative plus non-perturbative) contributions, i.e., $R(M^{2})\leq 0.05$. Moreover, the chosen range for $s_{0}$ should lie significantly above the production threshold of two $\Delta^{0}$ baryons, which is approximately $4m_{\Delta^{0}}^{2}\simeq 6.07~\mathrm{GeV}^{2}$. The range of $s_{0}$ may be selected as the minimal value that ensures a stable and reliable Borel window. Alternatively, $s_{0}$ can be fixed by fulfilling additional constraints so as to maximize the PC. Note that for multiquark hadrons, or more precisely exotic states, the available information is often scarce. In such situations, $s_{0}$ can be estimated based on the aforementioned constraints.
@@
It is important to emphasize that the present analysis excludes the influence of radiative corrections. Although the mass, extracted from the ratio of two sum rules, is relatively insensitive to such corrections, the residue calculated from an individual sum rule can exhibit considerable fluctuations. Furthermore, next-to-leading order (NLO) perturbative corrections to baryon sum rules have been reported as significant Krasnikov:1982ea ; Groote:1999zp . Similarly, while NLO $\alpha_{s}$ corrections to the perturbative component in the OPE for light tetraquark currents are sizable, their overall impact on the sum rules remains limited owing to the predominance of nonperturbative condensate contributions Groote:2014pva . Furthermore, given that the analysis of the $\Delta^{0}\Delta^{0}$ dibaryon with $J^{P}=2^{+}$ suggests the existence of a bound state, the estimation of its size can be interesting. Based on the definition of binding energy, the size of the $\Delta^{0}\Delta^{0}$ dibaryon can be determined using the following expression:
@@
where the binding energy is $B_{E}=2m_{\Delta^{0}}-m_{\Delta^{0}\Delta^{0}}$. In this expression, $r$ represents the distance between the components, while $\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}}$ corresponds to the reduced mass of the two-hadron system. The size of the tensor state is found to be approximately $r\simeq 0.90~\mathrm{fm}$, nearly twice that of the $\Delta^{0}\Delta^{0}$ dibaryon in the $J^{P}=0^{+}$ state. As is clear, both the tensor and scalar states remain below the confinement radius of 1 fm  Guo:2017jvc . In the case of the tensor state, while it may suggest a molecular configuration, the radius indicates an overlap between two $\Delta^{0}$ baryons, making the molecular description questionable. The obtained size suggests a compact hexaquark structure for this particle. However, calculating other properties of this state may help in achieving a better understanding of its internal configuration. In contrast to the $\Delta^{0}\Delta^{0}$ dibaryon, which represents a light-quark exotic dibaryon, it is of particular interest to estimate the radius of the $\Omega\Omega$ dibaryon mentioned above. Based on the binding energy calculated in Ref. Chen:2019vdh , and assuming the same quantum state as considered in our study, the resulting radii are found to be approximately $r\simeq 1.24~\mathrm{fm}$ for the $J^{P}=0^{+}$ state and $r\simeq 0.97~\mathrm{fm}$ for the $J^{P}=2^{+}$ state. These results indicate that the tensor state remains within the typical confinement radius.